OpenStudy (anonymous):

2x+6/5-x+7/3<3x-8.......SOLVE....

OpenStudy (jack1):

\[2x+6/5-x+7/3\] ?

OpenStudy (jack1):

or \[(2x+6)/(5-x)+7/3\]

OpenStudy (agent0smith):

You should put parentheses... this could be anything. I'll guess \[\Large \frac {2x+6} {5-x}+\frac{7} {3}< 3x-8\]

OpenStudy (anonymous):

\[\frac{ 2x+6 }{ 5 }-\frac{ x+7 }{3}<\frac{ 3x-8 }{ 2 }\]

OpenStudy (jack1):

you can re-write the LHS of the equation as: (x-17)/15 so now your equation looks something like: \[\frac{ x-17 }{ 15 }<\frac{ 3x-8 }{ 2 }\]

OpenStudy (jack1):

so if x = 3 or greater, the above holds true

OpenStudy (jack1):

if x = 2, LHS = RHS and if x = 1 or less LHS > RHS